# Center Biophysics and Biochemistry, (IVIC), Caracas, Venezuela

### v2

@@ Abstract
This type of noise in its white and Brownian version, also referred to as Brown noise, is presented in the Figure 8. Figures 8A and 8B, are white and brown noise with a pdf $\text{{{N}}}\left\{0.1\right\}$ and the Figures 8 and 8C with $\text{{{N}}}\left\{0.10^{-4}\right\}$. Note that if you look at the Figures 8A and 8C, or the Figures 8A and 8D without noticing the ordinate scale or variances noted above, one of those pairs looks the same as each other; the same goes for Figures 8B and 8C; These similarities are the so-called self-similarities or self affinities of fractal processes, which in the case of white and brown noises determine that they look the same at any scale of the abscissa that are observed. Note that despite the differences in their $g(x_{i}\;\mid\;\{\theta\})$ the figures 8B and 8D look very similar to each other and to Figures 6A and 6E.
The fractal dimension $D_{HB}=2$, for a white noise and is $D_{HB}=1.5$ for one brown [92, Sevcik1998a, Sevcik2010, Sevcik2017b]. Rodriguez-Hernandez and Sevcik [RodriguezHernandez2022] showed that $D_{S}\rightarrow 2$ for Figures 8A and 8C when the sequences lengthen towards $\aleph_{0}$ and that at the same time $D_{S}\rightarrow 1.5$ when the noise sequences Figures 8B and 8D are also lengthened. Rodríguez-Hernández and Sevcik [RodriguezHernandez2022] is therefore a demonstration of the existence, perhaps unsuspected, that in the operation of a modern hospital, there may be hidden factors, which make it behave like most service providers as a queue queue in contemporary literature in English), which in this case determines periods of low use of the Hospital, which can be considered “inefficiency” but which are not soluble if the hospital network of Madrid en bloc is not considered, not as individual hospitals of SERNAS. Please review the original work [RodriguezHernandez2022] for more details.
Figure 9: The relationship between open boxes and open circles to calculate the fractal dimension of a curve. If we call $\epsilon$ to the radius of the circle, then the side of the square will be $l=2\cdot\epsilon$. Please note that the lines have been drawn very thickly, they should be points of radius $d\rightarrow 0$ and the outermost layer of points, in a square or a circle “open”, are not included in the respective figure. From the figure it is intuitively obvious that, except for possible segments of a curve parallel to the abscissa or the ordinate of the curve, it will take fewer squares than circles to cover the curve since $d^{2}>\pi\epsilon^{2}=\pi\left(\tfrac{d}{2}\right)^{2}\impliedby 1>\tfrac{\pi}{4}$. Where $\epsilon$ is used as in Eq. (6) and subsequent ones.
@@
In the fractal literature [92] there are two ways to approximate the fractal dimension of a waveform. One of them is the use of open circles that cover the curve, this we have discussed in relation to $D_{S}$ [Sevcik1998a, Sevcik2010] where a double linear transformation to a unit square is used. The other is the use of “open boxes” that cover the wave, those boxes are usually squares of the do $l$. Long ignored by the authors, there is another form of fractal dimension of boxes, by Higuchi [96, 97]. It is so called because the waveform whose fractal dimension is to be estimated is covered by ce “boxes”, actually, in two dimensions, squares of side $d$. As follows from Figure 9 it takes more circles than squares to cover any waveform other than a vertical or horizontal line. In other circumstances $D_{\blacksquare}\lessapprox D_{\bullet}$.
Here we use a slightly different notation than Higuchi [96], using matrix notation (as in the rest of the book) and replace the use of “;” in favor of an equal sign between matrices. To understand the fractal dimension of Higuchi [96, 97] we will consider a time series of observations such as
@@
Here Higuchi introduces a condition that he does not explain, this is that the term in square brackets, there is none other than $\left[\tfrac{N-m}{k}\right]$, “denotes Gauss’s notation” (there is no other explanation about Gauss swimming that) and that both $k$ and $m$ are integers [we presume that it means $(m\in\mathbb{Z})\land(m\in\mathbb{Z})$] and points out that $m$ and $k$ indicate the initial time and the time interval (Between the points?) respectively. Starting from that series Higuchi [96] constructs the following matrix of equations, for say $N=100$:
$\begin{matrix}\{X_{1}^{3}\}&=&\left\{x(1),x(4),x(7),\ldots,x(97),x(100)\right\}\\ \{X_{2}^{3}\}&=&\left\{x(2),x(5),x(8),\ldots,x(95),x(98)\right\}\\ \{X_{3}^{3}\}&=&\left\{x(3),x(6),x(9),\ldots,x(96),x(99)\right\}\\ \end{matrix}$
@@
Eq. (26) in its first form is the Higuchi curve length equation with more contemporary notation, Eq. (27) is the same, simplified to make it more understandable and compact. If the original publication [96] is reviewed, there is no explanation about the two vertical lines that limit the term within the sum of the Eqs. (26) or (27), the journal where the article was published had no problem representing parentheses, braces or brackets of any size, as is obvious in Higuchi’s article [96] itself, so that these unexplained velicar lines can only be considered as indicating that the summation is made over the absolute values of the differences within the summation. This is strange, if $p_{i}$ and $p_{i+1}$ are point in a two dimensional Euclidean space with coordinates $(x_{i},y_{i})$ and $(x_{i+1},y_{i+1})$. respectively. The distance between $p_{i}$ and $p_{i+1}$, according to the Pitagoras theorem will be $\sqrt{(x_{i+1}-x_{i})^{2}+(y_{i+1}-y_{i})^{2}}$ and the absolute value function is never necessary.
According to Higuchi [96], the fraction $\frac{N-1}{[N-m]\cdot k}$ represents the “normalization of curve length” of a “time series subsystem”. Also, according to Higuchi also [96]: the length of the curve is defined for the time interval “$k$, $\;\left\langle L(k)\right\rangle$”, as the average value of the parameter $k$ of the sets $L_{m}(k)$. Again according to Higuchi [96]: yes
@@
the fractal dimension of the curve is $D_{Hig}$ [^4] The subscript Hig is ours.
@@
Higuchi [96] validated his method to determine the fractal dimension tests his method with “simulated data” to which he applies his method. First, he applies his technique to the simulated data $\{y_{i}\}_{i=1,2.....N}$ with fractal dimension $D=1.5$, $y_{i}$ is generated as
@@
where $z_{j}$ is is assumed Gaussian with mean 0 and variance 1 [96]. “The value of $1000$ was taken arbitrarily to eliminate the effect of arbitrarily sampling $z_{i}$ de $y_{1}$[96]. higuchi used the following values for the interval $k=1,2,3,4$, and $k=\left[2^{\tfrac{j-1}{4}}\right]_{(j=11,12,13\ldots)}$ for $k>4$, “where Gauss notation” (sic, I don’t know what Higuchi means), here Higuchi is inconsistent now uses ($k\in\mathbb{R}$), observe that $k=2^{2.5}=5.65685\ldots$, for example.
@@
which is a form of Eq. (28) changing the proportion to equality, which means that any difference produced by this change falls within Higuchi’s $D$. The $d$ is entered with the same sense as it has in the Figure 9. I introduce the implication because Higuchi simply calls ‘fractal’ his $D$ [96], and it is Mandelbrot [137, pg. 15 and Ch. 39] who invents the term “fractal”, and associates it with the Hausdorff-Besicovitch dimension [93, 27]. The “luck” referred to indicates the lack of association formal between $D_{Hid}$ and $D_{HB}$ that is not formally established anywhere. Higuchi [96] simply define: $D_{HB}=D_{Hig}$.
Higuchi [96, pg. 280] gives an example of his definition of length in relation to Burlaga and Klein´s description [36] of these authors’ fractality of the interplanetary magnetic field, as follows
@@
Another incomprehensible thing is why Higuchi uses the absolute value of the differences between neighboring points to calculate the length $L(k)$ of his curves. If these curves are constituted (away from Higuchi’s notation) by a set of data pairs $\{x_{i},y_{i}\}_{i=1,2,\ldots,N}$ then the length of the curve will be $\mathfrak{L}=\sum_{i=2}^{N}\left[(y_{i}-y_{i-1})^{2}+(x_{i}-x_{i-1})^{2}\right]^{0.5}$ according to the Pythagorean theorem $L=\sum_{i=2}^{N}\lvert y_{i}-y_{i-1}\rvert$, as Higuchi seems to use [96], so that $\mathfrak{L}>L$, which means that Higuchi [96] underestimates the kargos of the curves fractal dimension claims to estimate, unless the author uses the two vertical bars as something distinct, and unexplained, the absolute value of a difference.
@@
The logarithms of a long time, $\log_{b}\left\langle L(k)\right\rangle$ ($b$ is any basis), for a time series, $y_{i}$ with $N=217$, is plotted as a function of $\log_{b}L(k)$ in Figure 1 of Higuchi [96]. This appears to be a kind of Monte Carlo simulation of a Brownian noise like the one shown and discussed in Figure 8B and D, but does not give any information about the reliability of the generator of the distributed random variable such as $\text{{{N}}}\left\{0,\sigma^{2}\right\}$ and today would not be considered acceptable for publication [Sevcik1998a, Sevcik2010, Sevcik2016, RodriguezHernandez2022]. Higuchi [97] makes a second publication with his method, in it he cleans up algebra somewhat and makes it less baroque. But again they do what look like Monte Carlo-like simulations, it seems. They have been described for today’s requirements. Monte Carlo, it seems. They have been described for today’s requirements. After the work of 1990 [97], the use of the “fractal dimension of Higuchi” disappears from the literature (in my experience) until the year 2004 when it reappears, used mainly in medicine, without any additional formal analysis [2, 55, Klonowski2005, Spacic2011, Spasic2011a, 90, 1, 49, 125, 126, Shamsi2021, 9].
Thus, ways of estimating $D$ for curves in two-dimensional two-dimensional Euclidean spaces began to appear developed by Higuchi [96, 97] and by Katz [114]. Both of Higuchi’s papers were published in Physica D: Nonlinear Phenomena [96, 97], backed by another Japanese physicist.As we discussed earlier, the papers by Higuchi [96, 97] are strange. The first paper with a “muddy” algebra and many unexplained definitions. Higuchi’s method, then, disappears from the literature for 15 years and is cited again after 2004 [2, Klonowski2005, 90, 1, 49, 127, Shamsi2021, 9], some of these references include attempts to validate Higuchi´s dimension [127].
@@
Another interesting way for data analysis introduced in Rodriguez-Hernandez and Sevcik [RodriguezHernandez2022] is frequency analysis. Frequency analysis is not part of the analysis of chaotic or fractal systems, and also includes strictly statistical time series concepts such as autocorrelations and autocovariances. We will continue to use the work [RodriguezHernandez2022] as an example of this analysis.se can refer to [156, 25, Smith1997] among others.
A classic way to show periodicities of mathematical functions in the time domain is to transform them to the frequency domain of functions, which is a sum of sines and sews, using a rapid transformation of (FFT, which we will keep) [28, Welch1967]. The transformed values for each frequency ($f$) are “complex numbers” ($\mathbb{C}$), $c_{f}$ composed of two components: one of them QS some class sometimes called simply its “real component” ($\Re$) component such that $a_{f}\in\left(\mathbb{R}\right)$ and a called its component “imaginary” $\left(\Im\right)$ which is also ($b_{f}\sqrt{-1}$). A complex number ($c_{f}$) has the form
@@
where parameters indicate sets of numbers, and $\in$ reads “belongs to”. The “energy”, expressed as a stable nonzero curve, or as increases in sharp peaks or variations in amplitude (such as the peak in the Figure 10B, something we might informally call as its “weight”), contributed by each frequency component to the energy of the total signal in a complicated process that fluctuates and is expressed as
@@
Figure 10: Analysis of the false periodicity of Rws. Using spectral density analysis ($\boldsymbol{{\psi_{f_{j}}}}$). The panels are: Panel A- A perfectly periodic signal, the trigonometric function $\zeta(x)=\sin(x\tfrac{\pi}{128})$; Panel B- $\psi_{f_{j}}$ of the sine function displayed in panel A; Panel C- $\psi_{f_{j}}$ of white Gaussian noise shown in Figure 8A; Panel D- $\psi_{f_{j}}$ of Brown’s noise [92] Rw shown in Figure 8B; Panel E- $\psi_{f_{j}}$ of the sequence of hospitalized patients ($\boldsymbol{{InP}}$) that appears in the A Panels, up to E of Figure 6A. The ordinate of the panels from B to E, is the power dissipated at each frequency. The function shown is the displayed sine function presents the dissipated energy all its energy in a single peak $\psi_{j}$ (Panel B). The sequence for the number of hospitalized patients $SD{f_{j}}$ corresponds to a Gaussian Rw without any underlying periodicity. All graphics $SD$ were fitted a window of Hann [28, 91]. The rapid oscillations observed to the right from Panel B to D are artifacts due to the sampling of discrete points, in the signal to perform the Furrier transformation. Rw is used in the original work cited as random walk, more details in the text of this book and the original work cited [RodriguezHernandez2022]
@@
An intuitive way to understand this is to consider a perfectly periodic signal such as a sinusoidal trigonometric series such as a sine function $\zeta(x)=\sin(\tfrac{2x\pi}{T})$ where $\pi=3.1415926\ldots$ and $T$ is a constant (called su period) that has the same units (meters, seconds, grams, volts, or any other) that has $x$. An example is shown in Figure 10A. Figure 10B, presents the power spectrum of the Figure 10A consisting of a single peak $f=\tfrac{1}{T}$ frequency. Please note that calculating $\boldsymbol{{\psi_{f}}}$ the actual signal is sampled at intervals of $\Delta x$ (equivalent to multiplying by something called Dirac comb and the FFT also contains the frequency content of the Dirac comb produces the burst of high frequency vibrations of vibrations the right half of each stroke [28].
@@
The Figure 10C is a spectrum $\boldsymbol{{\psi_{f}}}$ of Gaussian white noise shown in Figure 8A; the ordering of panels from 8B to 8E, in arbitrary units. In the Figure 10C it is observed that all frequencies have the same ‘weight’, that is, all frequencies contribute equally. The Figure 10D is a spectrum Rw of a Brownian noise $\boldsymbol{{\psi_{f}}}$; On the logarithmic double coordinate scale, the spectral density line is a decaying line with a slope of $\tfrac{1}{f^{2}}$. The Figure 10E is $\boldsymbol{{\psi_{f}}}$ of patients hospitalized each day ($\boldsymbol{{InP}}$ in the figure published in English) is shown in the Figure 6A.
@@
Even though the simulated sequences in the Figure 10 are only $10000$ points and those in are only $1329$ points long, there is an obvious similarity between the Brownian noise $\boldsymbol{{\psi_{f}}}$ (Figure 10D), and $\boldsymbol{{\psi_{f}}}$ of the sequence of patients in the Hospital ($\boldsymbol{{Inp}}$) between May 1, 2013 and December 19, 2017 (Figure 10E). In cases $\boldsymbol{{\psi_{f}}}$ are characteristic of the Brownian noise Rw and none of these $\boldsymbol{{\psi_{f}}}$s have peaks that could indicate a suggestion of any periodic component which could some periodicity, despite short lapses or long Brownian lapses the Rws that could give this impression.
@@
Figure 10C is a Gaussian white noise $\psi_{f_{j}}$ calculated for the signal in the Figure 8A. In Figure 10C all frequencies have the same “weight”, the contribution of all frequencies is the same, “white” for this kind of noise. The Figure 10D shows $\psi_{f_{j}}$ of the Brownian noise Rw in the Figure 8B, it can be seen that with logarithmic double coordinates the spectral density follows a straight line that decays with a slope of $\tfrac{1}{f^{2}}$. Finally, also in Figure 10E, $\psi_{f_{j}}$ is shown which corresponds to the sequence of daily hospitalized patients Figure 6A. Except for the simulated sequences in the Figure 10 where $10^{4}$ points are simulated and for Hospital data where they are 1329 points in length, there is an obvious similarity between the Brownian noise $\psi_{f_{j}}$, and the $\psi_{f_{j}}$ of patients discharged from the Hospital between May 1, 2013 and December 19, 2017. In both cases $\psi_{f_{j}}$ are characteristic of a Rw and none of those $\psi_{f_{j}}$ has some suggestive peak of periodic components that could explain any periodicity, despite short periods in Rws that seem to suggest it.
@@
Since Mandelbrot introduced the concept of ‘fractal’ and published his “Fractal Geometry of Nature” [135, 137, 142] the notion of fractal dimension and fractal geometry have invaded virtually every corner of the economy [137, 140, Yuan2008, Yuan2009], medicine [169, 45, Raghavendra2009]. multiple corners of biology [92], epidemiology [148, Sevcik1998a, Sevcik2010], physics, service systems [Thomas2004, RodriguezHernandez2022], seismology [62, 40, Telesca2005, Telesca2006], the study of reversible reactions and enzymes [Savageau1995, Xu2007] and much more, including so-called multifrantal systems with more than one fractal component [120, Yuan2009, 112, Silva2009, 83, 122, 121].

### v3

@@ Abstract
This type of noise in its white and Brownian version, also referred to as Brown noise, is presented in the Figure 8. Figures 8A and 8B, are white and Brown noise with a pdf $\text{{{N}}}\left\{0.1\right\}$ and the Figures 8C and 8 with $\text{{{N}}}\left\{0.10^{-4}\right\}$. Note that if you look at the Figures 8A and 8C, or the Figures 8B and 8D without noticing the ordinate scale or variances noted above, one of those pairs looks the same as each other; the same goes for Figures 8B and 8C; These similarities are the so-called self-similarities or self affinities of fractal processes, which in the case of white and Brown noises determine that they look the same at any scale of the abscissa that are observed. Note that despite the differences in their $g(x_{i}\;\mid\;\{\theta\})$ figures 8B and 8D look very similar to each other and to Figures 6A and 6E.
The fractal dimension $D_{HB}=2$, for a white noise and is $D_{HB}=1.5$ for one Brown noise [94, Sevcik1998a, Sevcik2010, Sevcik2017b]. Rodriguez-Hernandez and Sevcik [RodriguezHernandez2022] showed that $D_{S}\rightarrow 2$ for Figures 8A and 8C when the sequences lengthen towards $\aleph_{0}$ and that at the same time $D_{S}\rightarrow 1.5$ when the noise sequences Figures 8B and 8D are also lengthened. Rodríguez-Hernández and Sevcik [RodriguezHernandez2022] is therefore a demonstration of the existence, perhaps unsuspected, that in the operation of a modern hospital, there may be hidden factors, which make it behave like most service providers as a queue in contemporary literature in English), which in this case determines periods of low use of the Hospital, which can be considered “inefficiency” but which are not soluble if the hospital network of Madrid en bloc is not considered, not as individual hospitals of SERNAS. Please review the original work [RodriguezHernandez2022] for more details.
Figure 9: The relationship between open boxes and open circles to calculate the fractal dimension of a curve. If we call $\epsilon$ to the radius of the circle, then the side of the square will be $d=2\cdot\epsilon$. Please note that the lines have been drawn very thickly, they should be points of radius $\epsilon=\frac{d}{2}\rightarrow 0$ and the outermost layer of points, in an “open” square or a circle, are not included in the respective square or a circle. From the figure it is intuitively obvious that, except for possible segments of a curve parallel to the abscissa or the ordinate of the curve, it will take fewer squares than circles to cover the curve since $d^{2}>\pi\epsilon^{2}=\pi\left(\tfrac{d}{2}\right)^{2}\impliedby 1>\tfrac{\pi}{4}$. Where $\epsilon$ is used as in Eq. (6) and subsequent ones.
@@
In the fractal literature [94] there are two ways to approximate the fractal dimension of a waveform. One of them is the use of open circles that cover the curve, this we have discussed in relation to $D_{S}$ [Sevcik1998a, Sevcik2010] where a double linear transformation to a unit square is used. The other is the use of “open boxes” that cover the wave, those boxes are usually squares of side $d$. Long ignored by the authors, there is another form of fractal dimension estimation ussing boxes, developed by by Higuchi [98, 99]. It is called fractal dimension because the waveform whose fractal dimension is to be estimated is covered by “boxes”, actually, in two dimensions, squares of side $d$. As follows from Figure 9 it takes more circles than squares to cover any waveform other than a vertical or horizontal line. In other circumstances $D_{\blacksquare}\lessapprox D_{\bullet}$. Here we use a slightly different notation than Higuchi [98], using matrix notation. To understand the fractal dimension of Higuchi [98, 99] we will consider a time series of observations such as
@@
Here Higuchi introduces a condition that he does not explain, this is that the term in square brackets, there is none other than $\left[\tfrac{N-m}{k}\right]$, “denotes Gauss’s notation” (there is no other explanation about Gauss suggesting that) and that both $k$ and $m$ are integers [we presume that it means $(m\in\mathbb{Z})\land(m\in\mathbb{Z})$] and points out that $m$ and $k$ indicate the initial time and the time interval (Between the points?) respectively. Starting from that series Higuchi [98] constructs the following matrix of equations, for say $N=100$:
$\begin{matrix}\{X_{1}^{3}\}&=&\left\{x(1),x(4),x(7),\ldots,x(97),x(100)\right\}\\ \{X_{2}^{3}\}&=&\left\{x(2),x(5),x(8),\ldots,x(95),x(98)\right\}\\ \{X_{3}^{3}\}&=&\left\{x(3),x(6),x(9),\ldots,x(96),x(99)\right\}\\ \end{matrix}$ (26)
@@
Eq. (27) in its first form is the Higuchi curve length equation with more contemporary notation, Eq. (28) is the same, simplified to make it more understandable and compact. If the original publication [98] is reviewed, there is no explanation about the two vertical lines that limit the term within the sum of the Eqs. (27) or (28), the journal where the article was published had no problem representing parentheses, braces or brackets of any size, as is obvious in Higuchi’s article [98] itself, so that these unexplained vertical lines can only be considered as indicating that the summation is made over the absolute values of the differences within the summation. This is strange, assuming $p_{i}$ and $p_{i+1}$ are point in a two dimensional Euclidean space with coordinates $(x_{i},y_{i})$ and $(x_{i+1},y_{i+1})$, respectively. The distance between $p_{i}$ and $p_{i+1}$, according to the Pitagoras theorem will be $\sqrt{(x_{i+1}-x_{i})^{2}+(y_{i+1}-y_{i})^{2}}$ and the absolute value function is never necessary.
According to Higuchi [98], the fraction $\frac{N-1}{[N-m]\cdot k}$ represents the “normalization of curve length” of a “time series subsystem”. Also, according to Higuchi [98]: the length of the curve is defined for the time interval “$k$, $\;\left\langle L(k)\right\rangle$”, as the average value of the parameter $k$ of the sets $L_{m}(k)$. Again according to Higuchi [98]: if
@@
the fractal dimension of the curve is $D_{Hig}$ [^5] The subscript Hig is mine.
@@
Higuchi [98] validated his method to determine the fractal dimension tests with “simulated data” to which he applied his method. First, he applies his technique to the simulated data $\{y_{i}\}_{i=1,2.....N}$ with fractal dimension $D=1.5$, $y_{i}$ is generated as
@@
where $z_{j}$ is is assumed Gaussian with mean 0 and variance 1 [98]. “The value of $1000$ was taken arbitrarily to eliminate the effect of arbitrarily sampling $z_{i}$ for $y_{1}$[98]. Higuchi used the following values for the interval $k=1,2,3,4$, and $k=\left(2^{\frac{j-1}{4}}\right)_{j=11,12,13\ldots}$ for $k>4$, “where Gauss notation” (sic, I don’t know what Higuchi means), here Higuchi is inconsistent now uses ($k\in\mathbb{R}$), observe that $k=2^{2.5}=5.65685\ldots$, for example.
@@
which is a form of Eq. (29) changing the proportion to equality, which means that any difference produced by this change falls within Higuchi’s $D$. The $d$ is entered with the same sense as it has in the Figure 9. I introduce the implication because Higuchi simply calls ‘fractal’ his $D$ [98], and it is Mandelbrot [139, pg. 15 and Ch. 39] who invents the term “fractal”, and associates it with the Hausdorff-Besicovitch dimension [95, 27]. The “luck” referred to indicates the lack of formal association between $D_{Hid}$ and $D_{HB}$ that is not formally established anywhere by Higuchi. Higuchi [98] simply defines: $D_{HB}=D_{Hig}$.
Higuchi [98, pg. 280] gives an example of his definition of length in relation to Burlaga and Klein´s description [37] of these authors’ fractality of the interplanetary magnetic field, as follows
@@
Another incomprehensible thing is why Higuchi uses the absolute value of the differences between neighboring points to calculate the length $L(k)$ of his curves. If these curves are constituted (away from Higuchi’s notation) by a set of data pairs $\{x_{i},y_{i}\}_{i=1,2,\ldots,N}$ then the length of the curve will be
$\mathfrak{L}=\sum_{i=2}^{N}\sqrt{(y_{i}-y_{i-1})^{2}+(x_{i}-x_{i-1})²}$
according to the Pythagorean theorem $L=\sum_{i=2}^{N}\lvert y_{i}-y_{i-1}\rvert$, as Higuchi seems to use [98], so that $\mathfrak{L}>L$, which means that Higuchi [98] underestimates the length of the curve’s fractal dimension it claims to estimate, unless the author uses the two vertical bars as something distinct, and unexplained.
@@
The logarithms of waveform’s length at a long time, $\log_{b}\left\langle L(k)\right\rangle$ ($b$ is any basis), for a time series, $y_{i}$ with $N=217$, is plotted as a function of $\log_{b}L(k)$ in Figure 1 of Higuchi [98]. This appears to be a kind of Monte Carlo simulation of a Brownian noise like the one shown and discussed in Figure 8B and 8D, but does not give any information about the reliability of the generator of a random variable such as $\text{{{N}}}\left\{0,\sigma^{2}\right\}$ and today would not be considered acceptable for publication [Sevcik1998a, Sevcik2010, Sevcik2016, RodriguezHernandez2022]. Higuchi [99] makes a second publication with his method, in it he cleans up algebra somewhat and makes it less baroque. But again the 1990 [99] publication he does what looks like a Monte Carlo simulations which is not described. They have been described for today’s requirements. Monte Carlo, it seems. After the work of 1990 [99], the use of the “fractal dimension of Higuchi” disappears from the literature (in my experience) until the year 2004 when it reappears, and is used used mainly in medicine, without any additional formal analysis [2, 56, Klonowski2005, Spacic2011, Spasic2011a, 92, 1, 50, 127, 128, Shamsi2021, 8].
Thus, ways of estimating $D$ for curves in two-dimensional Euclidean spaces began to appear developed by Higuchi [98, 99] and by Katz [116]. Both of Higuchi’s papers were published in Physica D: Nonlinear Phenomena [98, 99], backed by another Japanese physicist.As we discussed earlier, the papers by Higuchi [98, 99] are strange. The first paper with a “muddy” algebra and many unexplained definitions. Higuchi’s method, then, disappears from the literature for 15 years and is cited again after 2004 [2, Klonowski2005, 92, 1, 50, 129, Shamsi2021, 8], some of these references include attempts to validate Higuchi´s dimension [129].
@@
Another interesting way for data analysis of time series, used for example in Rodriguez-Hernandez and Sevcik [RodriguezHernandez2022], is frequency analysis. Frequency analysis is not part of the analysis of chaotic or fractal systems, and also includes strictly statistical time series concepts such as autocorrelations and autocovariances. We will continue to use the work [RodriguezHernandez2022] as an example of this analysis.se can refer to [158, 25, Smith1997] among others.
A classic way to show periodicities of mathematical functions in the time domain is to transform them to the frequency domain, which is a sum of sines and cosines, using a fast Fourier transformation (FFT) [28, Welch1967]. The transformed values for each frequency ($f$) are “complex numbers”, $c_{f}\in\mathbb{C}$, with two components: one of them is some class sometimes called simply “real component” such as $a_{f}\in\mathbb{R}$ and a $b_{f}\in\Im$ component called “imaginary”. A complex number has the form
@@
where parameters indicate sets of numbers, and $\in$ reads “belongs to”. The “energy”, expressed as a stable nonzero curve, or as increases in sharp peaks or variations in amplitude (such as the peak in the Figure 10B, this is called “spectral density” ($\boldsymbol{{\psi_{f}}}$) and, sometimes, we might informally call as the “weight”) , contributed by each frequency component to the energy of the total signal in a complicated process that fluctuates and is expressed as
@@
Figure 10: Analysis of the false periodicity of Rws. Using spectral density analysis ($\boldsymbol{{\psi_{f_{j}}}}$). The panels are: Panel A- A perfectly periodic signal, the trigonometric function $\zeta(x)=\sin(x\tfrac{\pi}{128})$; Panel B- $\psi_{f_{j}}$ of the sine function displayed in panel A; Panel C- $\psi_{f_{j}}$ of white Gaussian noise shown in Figure 8A; Panel D- $\psi_{f_{j}}$ of Brown’s noise [94] Rw shown in Figure 8B; Panel E- $\psi_{f_{j}}$ of the sequence of hospitalized patients ($\boldsymbol{{InP}}$) that appears in the Panels A up to E of Figure 6A. The ordinate of the panels from B to E, is the power dissipated at each frequency. The function shown is the displayed sine function presents the dissipated energy all its energy in a single peak $\psi_{j}$ (Panel B). The sequence for the number of hospitalized patients $SD{f_{j}}$ corresponds to a Gaussian Rw without any underlying periodicity. All graphics $SD$ were fitted a window of Hann [28, 93]. The rapid oscillations observed to the right from Panel B to D are artifacts due to the sampling of discrete points, in the signal to perform the Furrier transformation. Rw is used in the original work cited as random walk, more details in the text of this book and the original work cited [RodriguezHernandez2022]
@@
An intuitive way to understand this is to consider a perfectly periodic signal such as a sinusoidal trigonometric series such as a sine function $\zeta(x)=\sin(\tfrac{2x\pi}{T})$ where $\pi=3.1415926\ldots$ and $T$ is a constant (called period) that has the same units (meters, seconds, grams, volts, or any other) that has $x$. An example is shown in Figure 10A. Figure 10B, presents the power spectrum of Figure 10A consisting of a single peak with frequency $f=\tfrac{1}{T}$. Please note that calculating $\boldsymbol{{\psi_{f}}}$ the actual signal is sampled at intervals of $\Delta x$ (equivalent to multiplying by something called Dirac comb and the FFT also contains the frequency content of the Dirac comb produces the burst of high frequency vibrations of vibrations the right half of each stroke [28].
@@
Figure 10C is a spectrum $\boldsymbol{{\psi_{f}}}$ of Gaussian white noise shown in Figure 8A; the ordering of panels from 8B to 8E, in arbitrary units. In Figure 10C it is observed that all frequencies have the same “weight”, that is, all frequencies contribute equally. Figure 10D is a spectrum Rw of a Brownian noise $\boldsymbol{{\psi_{f}}}$; On the logarithmic double coordinate scale, the spectral density line is a decaying line with a slope of $\tfrac{1}{f^{2}}$. Figure 10E is $\boldsymbol{{\psi_{f}}}$ of patients hospitalized each day ($\boldsymbol{{InP}}$ in the figure) is shown in Figure 6A.
@@
Even though the simulated sequences in Figure 10 are only $10000$ points and those in are only $1329$ points long, there is an obvious similarity between the Brownian noise $\boldsymbol{{\psi_{f}}}$ (Figure 10D), and $\boldsymbol{{\psi_{f}}}$ of the sequence of patients in the Hospital ($\boldsymbol{{Inp}}$) between May 1, 2013 and December 19, 2017 (Figure 10E). In cases $\boldsymbol{{\psi_{f}}}$ are characteristic of the Brownian noise Rw and none of these $\boldsymbol{{\psi_{f}}}$s have peaks that could indicate a suggestion of any periodic component which could some periodicity, despite short lapses or long Brownian lapses the Rws that could give this impression.
@@
Figure 10C is a Gaussian white noise $\psi_{f_{j}}$ calculated for the signal in Figure 8A. In Figure 10C all frequencies have the same “weight”, the contribution of all frequencies is the same, “white” for this kind of noise. Figure 10D shows $\psi_{f_{j}}$ of the Brownian noise Rw in Figure 8B, it can be seen that with logarithmic double coordinates the spectral density follows a straight line that decays with a slope of $\tfrac{1}{f^{2}}$. Finally, also in Figure 10E, $\psi_{f_{j}}$ is shown which corresponds to the sequence of daily hospitalized patients Figure 6A. Except for the simulated sequences in Figure 10 where $10^{4}$ points are simulated and for Hospital data where they are 1329 points in length, there is an obvious similarity between the Brownian noise $\psi_{f_{j}}$, and the $\psi_{f_{j}}$ of patients discharged from the Hospital between May 1, 2013 and December 19, 2017. In both cases $\psi_{f_{j}}$ are characteristic of a Rw and none of those $\psi_{f_{j}}$ has some suggestive peak of periodic components that could explain any periodicity, despite short periods in Rws that seem to suggest it.
@@
Since Mandelbrot introduced the concept of ‘fractal’ and published his “Fractal Geometry of Nature” [137, 139, 144] the notion of fractal dimension and fractal geometry have invaded virtually every corner of the economy [139, 142, Yuan2008, Yuan2009], medicine [171, 46, Raghavendra2009]. Multiple corners of biology [94], epidemiology [150, Sevcik1998a, Sevcik2010], physics, service systems [Thomas2004, RodriguezHernandez2022], seismology [64, 41, Telesca2005, Telesca2006], the study of reversible reactions and enzymes [Savageau1995, Xu2007] and much more, including so called multifrantal systems with more than one fractal component [122, Yuan2009, 114, Silva2009, 85, 124, 123].